# Low-Dimensional Adaptation of Rectified Flow: A New Perspective through the Lens of Diffusion and Stochastic Localization

### v1

@@ Appendix C Proof of Lemmas related to Stochastic Localization (Section 3)
We begin with the SDE of $\tilde{Z}_{t}$:
@@
Now, we define $\overleftarrow{Y}_{\tau}:=\frac{\sqrt{\omega_{N-\tau}}\tilde{Z}_{t(N-\tau)}}{t(N-\tau)}$. We first begin by noting that $\frac{\sqrt{\omega_{N-\tau}}}{t(N-\tau)}=\frac{1}{\sigma_{t(N-\tau)}}$. We also have $\log\omega_{N-\tau}=-2\int_{0}^{N-\tau}\beta(u)du$. Therefore, we have
@@
$\Rightarrow\frac{d\tau}{dt}=\frac{1}{2\beta(N-\tau)\omega(N-\tau)}.\frac{d\omega_{N-\tau}}{dt}=\frac{1}{2\beta(N-\tau)\omega(N-\tau)}.\frac{d(t^{2}/\sigma_{t}^{2})}{dt}$
@@
Next, we define the backward time $t_{\tau}:=t(N-\tau)$, and then we can write $\tilde{Z}_{t_{\tau}}=\sigma_{t_{\tau}}\overleftarrow{Y}_{\tau}$. This also yields that the score of $\overleftarrow{Y}_{\tau}$ satisfies $s^{\leftarrow}_{\tau}(\overleftarrow{Y}_{\tau})=\sigma_{t_{\tau}}s_{t}(\tilde{Z}_{t_{\tau}})$. Moreover, (14) yields that $\sigma_{t_{\tau}}s_{t}(\tilde{Z}_{t_{\tau}})=\nabla\log q_{N-\tau}(\overleftarrow{Y}_{\tau})$. Next, with (53) and (51) in mind, we can write
@@
This shows that $\overleftarrow{Y}_{\tau}$ satisfies SDE (12).
@@ D.1 Controlling posterior covariance of RF
$=\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right)\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]$
@@
The inequality follows from $e^{x}-1\leq x$ for $x\in(0,1)$, and $\frac{\eta_{t_{i}}}{(1-t_{i})t_{i}}\leq 2h$. Substituting $t_{i-1}$ by $t_{i}$ in the above inequality, we get the following
@@
$\leq-\int_{t_{i}}^{t_{i+1}}\frac{2t}{3(1-t)^{3}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})dt+\frac{8h\eta_{t_{i}}}{(1-t_{i+1})^{2}N^{10}}$
$\leq-\frac{2\eta_{t_{i}}(t_{i+1}+t_{i})}{3(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})+\frac{8h\eta_{t_{i}}}{(1-t_{i+1})^{2}N^{10}}$
@@
$\frac{\eta_{t_{i}}t_{i}}{(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})\leq\frac{3}{4}\left(\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i}})]-\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i+1}})]\right)+\frac{6h\eta_{t_{i}}}{(1-t_{i+1})^{2}N^{10}}.$ (59)
D.3 Proof of Claim (38)

### v2

@@ Appendix C Proof of Lemmas related to Stochastic Localization (Section 3)
We begin with the solution $\tilde{Z}_{t}$ of STOC-RF SDE (6):
@@
Now, we define $\overleftarrow{Y}_{\tau}:=\frac{\sqrt{\omega_{N-\tau}}\tilde{Z}_{t(N-\tau)}}{t(N-\tau)}$. We first begin by noting that
$\frac{\sqrt{\omega_{N-\tau}}}{t(N-\tau)}=\frac{1}{\sigma_{t(N-\tau)}}$ (52)
We also have $\log\omega_{N-\tau}=-2\int_{0}^{N-\tau}\beta(u)du$. Therefore, we have
@@
$\Rightarrow\frac{d\tau}{dt}\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{2\beta(N-\tau)\omega(N-\tau)}.\frac{d\omega_{N-\tau}}{dt}=\frac{1}{2\beta(N-\tau)\omega(N-\tau)}.\frac{d(t^{2}/\sigma_{t}^{2})}{dt}$
@@
Step (i) follows from (52). Next, we define the backward time $t_{\tau}:=t(N-\tau)$, and then we can write $\tilde{Z}_{t_{\tau}}=\sigma_{t_{\tau}}\overleftarrow{Y}_{\tau}$. This also yields that the score of $\overleftarrow{Y}_{\tau}$ satisfies $s^{\leftarrow}_{\tau}(\overleftarrow{Y}_{\tau})=\sigma_{t_{\tau}}s_{t}(\tilde{Z}_{t_{\tau}})$. Moreover, (14) yields that $\sigma_{t_{\tau}}s_{t}(\tilde{Z}_{t_{\tau}})=\nabla\log q_{N-\tau}(\overleftarrow{Y}_{\tau})$. Next, with (54) and (51) in mind, we can write
@@
This shows that $\overleftarrow{Y}_{\tau}$ satisfies the backward DDPM SDE (12).
@@ D.1 Controlling posterior covariance of RF
Proof.
Now, we note that $\mathbf{A}_{s}$ and $\tilde{\boldsymbol{\Sigma}}_{t(s)}$ (so is ${\boldsymbol{\Sigma}}_{t(s)}$) are marginally equal in law. Therefore, by chain rule we have the following:
$\frac{d}{dt}\mathbb{E}({\boldsymbol{\Sigma}}_{t})$ $=\frac{d}{dt}\mathbb{E}(\tilde{\boldsymbol{\Sigma}}_{t})$
$=\frac{d}{dt}\mathbb{E}(\mathbf{A}_{s})$
$=\frac{2t}{(1-t)^{3}}\frac{d}{ds}\mathbb{E}(\mathbf{A}_{s})$
$=-\frac{2t}{(1-t)^{3}}\mathbb{E}(\mathbf{A}_{s}^{2})$
$=-\frac{2t}{(1-t)^{3}}\mathbb{E}({\boldsymbol{\Sigma}}_{t}^{2}).$
@@
$\leq\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right)\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]$
@@
The inequality follows from $e^{x}-1\leq x$ for $x\in(0,1)$, and $\frac{\eta_{i}}{(1-t_{i})t_{i}}\leq 2h$ (Lemma D.2). Substituting $t_{i-1}$ by $t_{i}$ in the above inequality, we get the following
@@
$\leq-\int_{t_{i}}^{t_{i+1}}\frac{2t}{3(1-t)^{3}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})dt+\frac{8h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}$
$\leq-\frac{2\eta_{t_{i}}(t_{i+1}+t_{i})}{3(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})+\frac{8h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}$
@@
$\frac{\eta_{t_{i}}t_{i}}{(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})\leq\frac{3}{4}\left(\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i}})]-\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i+1}})]\right)+\frac{6h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}.$ (60)
D.3 Properties of time scheduling (15)
Lemma D.2.
The time-steps (15) satisfies the following for $i\in\{1,\ldots,N-2\}$:
  1. (a)
    We have $\eta_{i}=ht_{i}{\mathbbm{1}}\left\{{t_{i}\leq 1/2}\right\}+h(1-t_{i+1}){\mathbbm{1}}\left\{{t_{i}>1/2}\right\}$.
  2. (b)
    We have $\frac{\eta_{i}}{1-t_{i}}\leq h$ and $\frac{\eta_{i}}{1-t_{i+1}}\leq h$.
  3. (c)
    We have $\frac{\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\leq h^{2}$, as $\frac{\eta_{i}}{(1-t_{i+1})}\leq h$ and $(1-t_{i})\leq 1$.
  4. (d)
    We have $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}=\sum_{i:0<t_{i}<1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}+\sum_{i:t_{i}\geq 1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}\leq 4h^{2}N$.
Proof.
Ghost line
  • Part (a): If $t_{i}<1/2$, we have $\eta_{i}=t_{i+1}-t_{i}=(1+h)t_{i}-t_{i}=ht_{i}$. For the case $t_{i}\geq 1/2$, we have
    $\eta_{i}$ $=t_{i+1}-t_{i}$
    $=(1-t_{i})-(1-t_{i+1})$
    $=h(1-t_{i+1}).$
  • Part (b): For $t_{i}<1/2$, we have $\frac{\eta_{i}}{1-t_{i}}=ht_{i}/(1-t_{i})\leq h$. We use the fact that $\frac{t}{1-t}<1$ for $t\leq 1/2$. Similarly, $\frac{\eta_{i}}{1-t_{i+1}}=\frac{ht_{i}}{1-t_{i+1}}$. Now, note that $t_{i}<1/2$ implies that $t_{i+1}\leq 1/2$. Therefore, $\frac{t_{i}}{1-t_{i+1}}\leq 1$, and it yields $\frac{\eta_{i}}{1-t_{i+1}}=\frac{ht_{i}}{1-t_{i+1}}<h$.
    For $t_{i}\geq 1/2$, we have $\frac{\eta_{i}}{1-t_{i}}=\frac{h(1-t_{i+1})}{1-t_{i}}<h$. Similarly, $\frac{\eta_{i}}{1-t_{i+1}}=\frac{h(1-t_{i+1})}{1-t_{i+1}}=h$. This yields the coveted bound.
  • Part (c): The proof immediately follows from Part (b).
  • Part (d): Fist, from Part (a), we can note that $\frac{\eta_{i}}{t_{i}}\leq h$. Also, Part (b) yields that $\frac{\eta_{i}}{1-t_{i+1}}\leq h$. Therefore,
    $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}$
    $=\sum_{i:0<t_{i}<1/2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\eta_{i}^{2}}{t_{i}^{2}}}.\frac{1}{(1-t_{i+1})^{2}}+\sum_{i:t_{i}\geq 1/2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}}}.\frac{1}{t_{i}^{2}}$
    $\leq\sum_{i:0<t_{i}<1/2}4h^{2}+\sum_{i:t_{i}\geq 1/2}4h^{2}$
    $\leq 4h^{2}N.$